Circular No. 8098

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 F1 (LINEAR) An apparently asteroidal object reported by the LINEAR project (dis-

covery observation below), and posted on the NEO Confirmation Page, has been found to have cometary appearance by numerous CCD observers on Mar. 24–25 (who have reported diffuseness, a coma diameter of up to 10′′, and a tail of length 7′′–40′′ in p.a. 300^{o}–330^{o}), including A. R. Apitzsch (Wildberg, Germany); S. Sanchez, R. Stoss, and J. Nomen (Mallorca, Spain); M. Fröhlich (Essen, Germany); G. Hug (Eskridge, KS; m_{1} = 16.7); J. E. Arlot (Observatoire de Haute Provence); H. Mikuž (Crni Vrh, Slovenia); and L. Buzzi (Varese, Italy; m_{1} = 17.7). Available astrometry, very preliminary parabolic orbital elements (q = 3.9 AU, T = 2002 July 31), and an ephemeris appear on MPEC 2003-F35.

2003 UT  α_{2000}  δ_{2000}  m_{2}
Mar. 23.43063  17^{h}23^{m}30.^{s}78  +51^{o}22′52′′.1  18.2

SUPERNOVAE 2003bn, 2003bu, 2003cb, 2003ch, AND 2003ci M. Salvo and B. Schmidt, Research School of Astronomy and Astro-

physics, Australian National University (ANU); and J. Tonry, Institute for Astronomy, University of Hawaii, report that a spectrum (range 350–900 nm) of SN 2003bn (cf. IAUC 8088), obtained on Mar. 24.48 UT with the ANU 2.3-m telescope (+ DBS spectrograph), shows it to be a type-II supernova, ≈ 1 month past explosion. The redshift of the host galaxy is cz = 3800 km/s as determined from narrow Hα at the position of the supernova. Weak iron lines in the supernova spectrum indicate a photospheric velocity of 5400 km/s. A spectrum of SN 2003bu (cf. IAUC 8092), obtained on Mar. 24.68, shows it to be a type-Ic supernova; cross correlation with a library of type-Ic supernovae spectra indicates that SN 2003bu is ≈ 1 week past maximum light. A spectrum of SN 2003cb (cf. IAUC 8094), obtained on Mar. 24.64, shows it to be a type-II supernova, ≈ 2 months after explosion; weak lines in the supernova spectrum indicate a photospheric velocity of 3000 km/s. A spectrum of SN 2003ch (cf. IAUC 8097), obtained on Mar. 24.45, shows it to be a type-Ia supernova. Cross correlation with a library of type-Ia supernovae spectra indicates that SN 2003ch is ≈ 3 ± 2 days before maximum light, and at a redshift of z = 0.025. A spectrum of SN 2003ci (cf. IAUC 8097), obtained on Mar. 25.52, shows it to be a type-II supernova, ≈ 3 weeks past explosion; weak iron lines in the spectrum indicate a photospheric velocity of 6500 km/s.

2003 March 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green